OpenStudy (anonymous):

how to solve x over 9 minus 1 equals negative 10

OpenStudy (anonymous):

Solve for x: x/9-1 = -10 Put the fractions in x/9-1 over a common denominator. Put each term in x/9-1 over the common denominator 9: x/9-1 = x/9-(9)/9: x/9-(9)/9 = -10 Combine x/9-(9)/9 into a single fraction. x/9-(9)/9 = (x-9)/9: (x-9)/9 = -10 Multiply both sides by a constant to simplify the equation. Multiply both sides of (x-9)/9 = -10 by 9: (9 (x-9))/9 = -10×9 Cancel common terms in the numerator and denominator of (9 (x-9))/9. (9 (x-9))/9 = 9/9×(x-9) = x-9: x-9 = -10×9 Multiply 9 and -10 together. 9 (-10) = -90: x-9 = -90 Isolate terms with x to the left hand side. Add 9 to both sides: x+(9-9) = 9-90 Look for two terms that sum to zero. 9-9 = 0: x = 9-90 Evaluate 9-90. 9-90 = -81: Answer: | | x = -81